# Gradient Flow Sampler-based Distributionally Robust Optimization

### v1

@@ 5 Optimization Complexity Analysis
The stochastic gradient estimator $\widehat{g}:=\mathbb{E}_{\widehat{\rho}_{Y|X=x}}[\nabla_{\theta}\ell(\theta,y)]$ has bounded variance $\sigma^{2}$. The inner-loop sampler (Algorithm 1) generates a distribution $\widehat{\rho}_{Y|X=x}$ with a uniformly bounded expected error $\mathbb{E}_{x\sim\widehat{\rho_{N}}}[W_{2}(\widehat{\rho}_{Y|X=x},\rho^{*}_{Y|X=x})]\leq\delta_{\text{sample}}$.
@@
For any fixed $\theta$ and $x$, the conditional distribution $\rho_{Y|X=x}$ is $L_{U}$-smooth and satisfies the $\alpha_{U}$-log-Sobolev inequality (LSI).
Assumption 4 is motivated by the convergence analysis of ULA (Vempala and Wibisono,, 2019; Wibisono,, 2025), which often requires the target distribution to satisfy the $\alpha$-log-Sobolev inequality ($\alpha$-LSI) (Gross,, 1975). We recall that a distribution $\rho$ satisfies the $\alpha$-LSI for some $\alpha>0$ if, for any probability distribution $\nu$, the following inequality holds:
$\text{FI}(\nu\|\rho)\geq 2\alpha\operatorname{KL}(\nu\|\rho),$
where $\operatorname{KL}(\nu\|\rho):=\mathbb{E}_{\nu}\left[\log\frac{\nu}{\rho}\right]$ is the KL divergence and $\text{FI}(\nu\|\rho):=\mathbb{E}_{\nu}\left[\left\|\nabla\log\frac{\nu}{\rho}\right\|^{2}\right]$ is the relative Fisher information. Since the relative Fisher information is the squared gradient norm of the KL divergence in the $W_{2}$ metric, the property that $\rho$ satisfies LSI is equivalent to the Polyak-Łojaciewicz (PL) or gradient domination property of the KL divergence objective function.
It is important to clarify, however, that the PL inequality for the potential function $\widetilde{V}_{x,\tau}$ itself (as in (18)) is not equivalent to the LSI for the distribution $\rho_{Y|X=x}=\exp\left(-\frac{2\tau}{\epsilon}\widetilde{V}_{x,\tau}\right)$.
Assumption 4 allows us to bound the number of inner ULA iterations required to achieve the necessary sampling accuracy. Combining these results yields the total complexity.
@@
$\text{Complexity}=O\left(\frac{L_{\Phi}L_{U}^{2}L_{f}^{2}d^{2}}{\alpha_{U}^{3}\epsilon_{\text{opt}}^{4}}\cdot\log\frac{1}{\epsilon_{\text{opt}}}\right)=\tilde{O}\left(\frac{L_{\Phi}L_{U}^{2}L_{f}^{2}d^{2}}{\alpha_{U}^{3}\epsilon_{\text{opt}}^{4}}\right).$
The detailed derivations for the gradient bias, the proof of Theorem 1, and the complexity analysis are provided in Appendix C.
@@ 6.1 Classification under Data Imbalance
Figure 1: Robust Decision Boundaries with Biased Data. (a) Decision boundaries learned by different methods on the biased circle dataset. The training data are shown as orange (positive class) and blue (negative class) points. The final classification boundaries are shown for the WRM (green), WGF (purple), WFR (red), and Dual (brown) models. All models were trained for 40 epochs. We set the regularization parameter $\tau=2.5$ for all methods and the entropy regularization $\epsilon=0.15$ for methods based on entropy-regularized Wasserstein DRO problem. (b) Samples from the worst-case distribution generated by WFR sampler at the first epoch. (c) Worst-case samples generated by WRM method at the first epoch. WRM can only generate discrete distributions as worst-case distribution while Sinkhorn DRO uses potentially continuous distributions as worst-case distribution. The original data points are shown as circles and the worst-case samples are shown by circles with a shallower color.
@@ 6.2 Two Moon Classification
To evaluate the capacity of generating the worst-case distribution of WRM, WGF, WFR, SVG, and RGO, we conducted an experiment on a pre-trained SAA model. This setup isolates the inner-loop optimization process used to find the worst-case distribution. The evolution of the inner objective function, $\mathbb{E}[\widetilde{V}_{x,\tau}(z)]$, is depicted in Figure 3.
@@ 7 Other Works and Discussion
Recent works have advanced the understanding and application of Distributionally Robust Optimization (DRO) in continuous probability spaces. (Xu et al.,, 2024) proposed FlowDRO, a practical map-based approach that learns a deterministic transport map, parameterized by a sequence of neural networks, to solve the standard Wasserstein DRO problem. (Zhu and Xie,, 2024) provide a foundational theoretical analysis, establishing convergence guarantees for a general iterative min-max framework for discrepancy-based DRO. Our work is rather orthognal to those efforts in that it addresses the DRO problems with a sample-based methodology that offers several aforementioned advantages. An important contribution of this paper is allowing optimizers to go beyond the standard (Wasserstein) gradient descent type of algorithm, inspiring novel approaches such SVG and WFR.
@@ A.1 Dual approach to Entropy-regularized Wasserstein DRO
The Sinkhorn DRO problem, as formulated by (Wang et al.,, 2021), aims to find a decision that minimizes the worst-case expected loss over an ambiguity set defined by Sinkhorn divergence, which is identical to entropy-regularized Wasserstein distance in this problem. Given a loss function $\ell(\theta,z)$, a nominal distribution $\widehat{\rho_{N}}$, and a radius $r>0$, the primal inner problem is:
@@
A dual approach, established by (Wang et al.,, 2021), reformulates the inner maximization problem. Using the strong duality of this problem and substituting (10) into this problem, the inner problem can be expressed as a minimization over a dual variable $\tau$:
@@
where $\rho_{x,\epsilon}(y)=\frac{\exp(\frac{-c(x,y)}{\epsilon})}{Z}$ is a kernel distribution centered at $x$, and $Z$ is a normalizing constant.
To approximate the value and the gradient of the nested expectation in (21), (Wang et al.,, 2021) employ a two-level sampling procedure for a given $\tau$: first, a sample $x$ is drawn from the nominal distribution $\widehat{\rho_{N}}$; second, a set of samples are drawn from the kernel distribution $\rho_{x,\epsilon}$ to approximate the inner expectation.
@@
Let us define a reference distribution $\rho_{0}(y)$ as a kernel density estimator of the empirical distribution $\widehat{\rho_{N}}=\frac{1}{N}\Sigma_{i=1}^{N}\delta(x_{i})$: $\rho_{0}(y)=\frac{1}{N}\Sigma_{i=1}^{N}K_{\epsilon}(y,x_{i})$, where $K_{\epsilon}(y,x_{i})=\alpha_{\epsilon}\exp(-c(y,x_{i})/\epsilon)$ and $\alpha_{\epsilon}=\mathbb{E}[\exp(-c(y,x_{i})/\epsilon)]^{-1}$. With this choice, and by setting $\tau=\tau^{\prime}/\epsilon$, the solution to the inner maximization problem is given by (Hu and Hong,, 2013):
@@
However, we must note that there is no global PL or strong convexity analogous to the case of the Wasserstein gradient flow, as discussed in the main text. A “warm-start” condition is necessary as it has been shown that global PL/gradient dominance cannot hold for the Fisher-Rao/(spherical)Hellinger gradient flows; see (Mielke and Zhu,, 2025; Carrillo et al.,, 2024) for details. There exist fine-grained warm-start initialization conditions in the gradient flow literature; see (Lu et al.,, 2023, 2019; Chen et al.,, 2023).
For example, (Lu et al.,, 2023) showed that, after a warm-start initialization in the form of a density-ratio lower bound, it is possible to get an improved exponential convergence rate than the pure Wasserstein gradient flow as characterized in Lemma 4. However, it is important to note that the specific warm-start initialization condition cannot be verified in our DRO problems. Nonetheless, in practice, we observe significant speedups using the WFR gradient flow as compared to the pure Wasserstein gradient flow.
@@ B.1 Using SVGD to Sample from the Worst-Case Distribution
7:   $\phi^{s,i}_{t}=\frac{1}{m}\sum_{j=1}^{m}\left[k(y^{s,i}_{t},y^{s,j}_{t})\nabla\widetilde{V}_{x^{s},\tau}(y^{s,j}_{t})+\nabla_{2}k(y^{s,i}_{t},y^{s,j}_{t})\right]$
8:   $y^{s,i}_{t+1}=y^{s,i}_{t}+\phi^{s,i}_{t}$
@@ B.2 Using RGO to Sample form the Worst-case Distribuion
which precisely matches the form of the problem (25) solved by backward step. This structural equivalence suggests that methods developed for proximal sampler can be directly applied to solve the entropy-regularized DRO problem. Specifically, we can employ the Restricted Gaussian Oracle (RGO) from (Lee et al.,, 2021) to solve this. The sampling process for a given $x\sim\widehat{\rho_{N}}$ involves two steps:
@@ C.4 Derivation of ULA Sampler Complexity (for Theorem 2)
The complexity of the ULA sampler depends on the geometric properties of the inner target distribution $\rho_{Y|X=x}$. Under Assumption 4, standard results on ULA convergence (Vempala and Wibisono,, 2019) state that after $T$ steps, the KL-divergence to the target is bounded. To achieve a final KL-divergence of $\delta_{KL}$, the number of iterations required is $T=O\left(\frac{L_{U}^{2}d}{\alpha_{U}^{2}\delta_{KL}}\log\frac{1}{\delta_{KL}}\right)=\tilde{O}\left(\frac{L_{U}^{2}d}{\alpha_{U}^{2}\delta_{KL}}\right)$.
@@
$W_{2}^{2}(\widehat{\rho}_{Y|X=x},\rho_{Y|X=x})\leq\frac{2}{\alpha_{U}}\operatorname{KL}(\widehat{\rho}_{Y|X=x}||\rho^{*}_{Y|X=x})$
@@
$\delta_{KL}\leq\frac{\alpha_{U}}{2}W_{2}^{2}(\widehat{\rho}^{*}_{Y|X=x},\rho_{Y|X=x})=O\left(\alpha_{U}\frac{\epsilon_{\text{opt}}^{2}}{L_{f}^{2}}\right).$
@@
$T_{ULA}=\tilde{O}\left(\frac{L_{U}^{2}d}{\alpha_{U}^{2}\cdot\alpha_{U}\frac{\epsilon_{\text{opt}}^{2}}{L_{f}^{2}}}\right)=\tilde{O}\left(\frac{L_{U}^{2}L_{f}^{2}d}{\alpha_{U}^{3}\epsilon_{\text{opt}}^{2}}\right).$
@@
$=O\left(\frac{1}{\epsilon_{\text{opt}}^{2}}\right)\times\tilde{O}\left(\frac{L_{U}^{2}L_{f}^{2}d}{\alpha_{U}^{3}\epsilon_{\text{opt}}^{2}}\right)\times O(d)$
$=\tilde{O}\left(\frac{L_{U}^{2}L_{f}^{2}d^{2}}{\alpha_{U}^{3}\epsilon_{\text{opt}}^{4}}\right).$

### v2

@@ 5 Optimization Complexity Analysis
The stochastic gradient estimator $\widehat{g}:=\mathbb{E}_{\widehat{\rho}_{Y|X=x}}[\nabla_{\theta}\ell(\theta,y)]$ has bounded variance $\sigma^{2}$, and $\sigma\leq\epsilon_{\text{opt}}$. The inner-loop sampler (Algorithm 1) generates a distribution $\widehat{\rho}_{Y|X=x}$ with a uniformly bounded expected error $\mathbb{E}_{x\sim\widehat{\rho_{N}}}[W_{2}(\widehat{\rho}_{Y|X=x},\rho^{*}_{Y|X=x})]\leq\delta_{\text{sample}}$.
@@
For any fixed $\theta$ and $x$, the conditional distribution $\rho_{Y|X=x}$ is $L_{U}$-smooth and satisfies the $\lambda_{U}$-log-Sobolev inequality (LSI); see (20).
[…]
@@
$\text{Complexity}=O\left(\frac{L_{\Phi}L_{U}^{2}L_{f}^{2}d^{2}}{\lambda_{U}^{3}\epsilon_{\text{opt}}^{4}}\cdot\log\frac{1}{\epsilon_{\text{opt}}}\right)=\tilde{O}\left(\frac{L_{\Phi}L_{U}^{2}L_{f}^{2}d^{2}}{\lambda_{U}^{3}\epsilon_{\text{opt}}^{4}}\right).$
The detailed derivations for the gradient bias, the proof of Theorem 1 and Theorem 2 are provided in Appendix C.
@@ 6.1 Classification under Data Imbalance
Figure 1: Robust Decision Boundaries with Biased Data. (a) Decision boundaries learned by different methods on the biased circle dataset. The training data are shown as orange (positive class) and blue (negative class) points. The final classification boundaries are shown for the WRM (green), WGF (purple), WFR (red), and Dual (brown) models. All models were trained for 40 epochs. We set the regularization parameter $\tau=2.5$ for all methods and the entropy regularization $\epsilon=0.15$ for methods based on entropy-regularized Wasserstein DRO problem. (b) Samples from the worst-case distribution generated by WFR sampler at the first epoch. (c) Worst-case samples generated by WRM method at the first epoch. WRM can only generate discrete distributions as worst-case distribution while entropy-regularized DRO uses potentially continuous distributions as worst-case distribution. The original data points are shown as circles with black edge and the worst-case samples are shown by circles with a shallower color.
@@ 6.2 Two Moon Classification
To evaluate the capacity of generating the worst-case distribution of WRM, WGF, WFR, SVG, and RGO, we conducted an experiment on a pre-trained SAA model. This setup isolates the inner-loop optimization process used to find the worst-case distribution. The evolution of the inner objective function, $\mathbb{E}[-\widetilde{V}_{x,\tau}(z)]$, is depicted in Figure 3.
@@ 7 Other Works and Discussion
Recent works have advanced the understanding and application of Distributionally Robust Optimization (DRO) in continuous probability spaces. Xu et al., (2024) proposed FlowDRO, a practical map-based approach that learns a deterministic transport map, parameterized by a sequence of neural networks, to solve the standard Wasserstein DRO problem. Zhu and Xie, (2024) provide a foundational theoretical analysis, establishing convergence guarantees for a general iterative min-max framework for discrepancy-based DRO. Our work is rather orthognal to those efforts in that it addresses the DRO problems with a sample-based methodology that offers several aforementioned advantages. An important contribution of this paper is allowing optimizers to go beyond the standard (Wasserstein) gradient descent type of algorithm, inspiring novel approaches such SVG and WFR.
@@ A.1 Dual approach to Entropy-regularized Wasserstein DRO
The Sinkhorn DRO problem, as formulated by Wang et al., (2021), aims to find a decision that minimizes the worst-case expected loss over an ambiguity set defined by Sinkhorn divergence, which is identical to entropy-regularized Wasserstein distance in this problem. Given a loss function $\ell(\theta,z)$, a nominal distribution $\widehat{\rho_{N}}$, and a radius $r>0$, the primal inner problem is:
@@
A dual approach, established by Wang et al., (2021), reformulates the inner maximization problem. Using the strong duality of this problem and substituting (10) into this problem, the inner problem can be expressed as a minimization over a dual variable $\tau$:
@@
where $\rho_{x,\epsilon}(y)=\frac{\exp(\frac{-c(x,y)}{\epsilon})}{Z}$ is a kernel distribution centered at $x$, and $Z$ is a normalization constant.
To approximate the value and the gradient of the nested expectation in (22), Wang et al., (2021) employ a two-level sampling procedure for a given $\tau$: first, a sample $x$ is drawn from the nominal distribution $\widehat{\rho_{N}}$; second, a set of samples are drawn from the kernel distribution $\rho_{x,\epsilon}$ to approximate the inner expectation.
@@
Let us define a reference distribution $\rho_{0}(y)$ as a kernel density estimator of the empirical distribution $\widehat{\rho_{N}}=\frac{1}{N}\Sigma_{i=1}^{N}\delta(x_{i})$: $\rho_{0}(y)=\frac{1}{N}\Sigma_{i=1}^{N}K_{\epsilon}(y,x_{i})$, where $K_{\epsilon}(y,x_{i})=\alpha_{\epsilon}\exp(-c(y,x_{i})/\epsilon)$ and $\alpha_{\epsilon}=\mathbb{E}[\exp(-c(y,x_{i})/\epsilon)]^{-1}$. With this choice, and by setting $\tau=\tau^{\prime}/\epsilon$, the solution to the inner maximization problem is given by Hu and Hong, (2013):
@@
However, we must note that there is no global PL or strong convexity analogous to the case of the Wasserstein gradient flow, as discussed in the main text. A “warm-start” condition is necessary as it has been shown that global PL/gradient dominance cannot hold for the Fisher-Rao/(spherical)Hellinger gradient flows; see Mielke and Zhu, (2025); Carrillo et al., (2024) for details. There exist fine-grained warm-start initialization conditions in the gradient flow literature; see Lu et al., (2023, 2019); Chen et al., (2023).
For example, Lu et al., (2023) showed that, after a warm-start initialization in the form of a density-ratio lower bound, it is possible to get an improved exponential convergence rate than the pure Wasserstein gradient flow as characterized in Lemma 4. However, it is important to note that the specific warm-start initialization condition cannot be verified in our DRO problems. Nonetheless, in practice, we observe significant speedups using the WFR gradient flow as compared to the pure Wasserstein gradient flow.
@@ B.1 Using SVGD to Sample from the Worst-Case Distribution
7:   $\phi^{s,i}_{t}=\frac{1}{m}\sum_{j=1}^{m}\left[-k(y^{s,i}_{t},y^{s,j}_{t})\cdot\frac{2\tau}{\epsilon}\nabla\widetilde{V}_{x^{s},\tau}(y^{s,j}_{t})+\nabla_{2}k(y^{s,i}_{t},y^{s,j}_{t})\right]$
8:   $y^{s,i}_{t+1}=y^{s,i}_{t}+\eta\phi^{s,i}_{t}$
@@ B.2 Using RGO to Sample form the Worst-case Distribuion
which precisely matches the form of the problem (26) solved by backward step. This structural equivalence suggests that methods developed for proximal sampler can be directly applied to solve the entropy-regularized DRO problem. Specifically, we can employ the Restricted Gaussian Oracle (RGO) from Lee et al., (2021) to solve this. The sampling process for a given $x\sim\widehat{\rho_{N}}$ involves two steps:
@@ C.3 Proof of Theorem 1 (Outer Loop Convergence)
Therefore, when $\delta_{\text{sample}}$ is controlled such that $\delta_{\text{sample}}=O(\epsilon_{\text{opt}}/L_{f})$, step size $r=O(1/L_{\Phi})$ and the iteration $S=O(1/\epsilon_{\text{opt}}^{2})$, $\frac{1}{S}\sum_{s=0}^{S-1}\mathbb{E}[\|\nabla\Phi(\theta^{s})|^{2}]\leq O(\epsilon_{\text{opt}}^{2})$.
@@ C.4 Derivation of ULA Sampler Complexity (for Theorem 2)
The complexity of the ULA sampler depends on the geometric properties of the inner target distribution $\rho_{Y|X=x}$. Under Assumption 4, standard results on ULA convergence (Vempala and Wibisono,, 2019) state that after $T$ steps, the KL-divergence to the target is bounded. To achieve a final KL-divergence of $\delta_{KL}$, the number of iterations required is $T=O\left(\frac{L_{U}^{2}d}{\lambda_{U}^{2}\delta_{KL}}\log\frac{1}{\delta_{KL}}\right)=\tilde{O}\left(\frac{L_{U}^{2}d}{\lambda_{U}^{2}\delta_{KL}}\right)$.
@@
$W_{2}^{2}(\widehat{\rho}_{Y|X=x},\rho_{Y|X=x})\leq\frac{2}{\lambda_{U}}\operatorname{KL}(\widehat{\rho}_{Y|X=x}||\rho^{*}_{Y|X=x})$
@@
$\delta_{KL}\leq\frac{\lambda_{U}}{2}W_{2}^{2}(\widehat{\rho}^{*}_{Y|X=x},\rho_{Y|X=x})=O\left(\lambda_{U}\frac{\epsilon_{\text{opt}}^{2}}{L_{f}^{2}}\right).$
@@
$T_{ULA}=\tilde{O}\left(\frac{L_{U}^{2}d}{\lambda_{U}^{2}\cdot\lambda_{U}\frac{\epsilon_{\text{opt}}^{2}}{L_{f}^{2}}}\right)=\tilde{O}\left(\frac{L_{U}^{2}L_{f}^{2}d}{\lambda_{U}^{3}\epsilon_{\text{opt}}^{2}}\right).$
@@
$=O\left(\frac{1}{\epsilon_{\text{opt}}^{2}}\right)\times\tilde{O}\left(\frac{L_{U}^{2}L_{f}^{2}d}{\lambda_{U}^{3}\epsilon_{\text{opt}}^{2}}\right)\times O(d)$
$=\tilde{O}\left(\frac{L_{U}^{2}L_{f}^{2}d^{2}}{\lambda_{U}^{3}\epsilon_{\text{opt}}^{4}}\right).$